James needs to create a nitric acid solution which is less than 15% nitric acid. He begins with 12 fluid ounces of a 20% nitric acid solution in water and he adds 8 fluid ounces of water. Does he need to add more water?
Answer by jim_thompson5910(35256)

"He begins with 12 fluid ounces of a 20% nitric acid solution in water". He has 12*0.20 = 2.4 fluid ounces of pure acid.


"he adds 8 fluid ounces of water" which means the 12 fluid ounces of solution turns into 12+8 = 20 fl. oz of solution


To find the new acid concentration, we divide the amount of pure acid by the amount of total solution: 2.4/20 = 0.12 = 12%


The new concentration is 12% which is less than 15%. He does not need to add more water.